OpenStudy (jadedry):

Solve the following equation:

OpenStudy (jadedry):

\[3^{2x+1} -26(3^x) - 9 = 0\] I have no idea where to start, could you please provide some hints? Thanks in advance!

OpenStudy (anonymous):

do u want answer in terms of x or want to simplify it?

ganeshie8 (ganeshie8):

Let \(3^x = t\)

OpenStudy (anonymous):

as ganeshie said but before line up the variables so everything clear like this

OpenStudy (anonymous):

\[-9-26 \times 3^{x}+3^{2x+1}=0\]

OpenStudy (anonymous):

\[-9-26 \times 3^{x}+3(3^{x})^{2}\]

OpenStudy (anonymous):

\[3t ^{2}-26t-9=0\]

OpenStudy (anonymous):

then factorise it

OpenStudy (jadedry):

@ayeshaafzal221 I believe I'm supposed to solve for X, I got the same equation you did, but my textbook says that the answer is 2?

OpenStudy (anonymous):

yes the answer is 2

OpenStudy (anonymous):

This is why math makes no sense... I read the equation, i caught on for the next part, then all that sht fell apart in my head

OpenStudy (anonymous):

\[(t-9)(3t+1)=0\]

OpenStudy (anonymous):

split this into two eq

OpenStudy (anonymous):

\[t-9=0 , 3y+1=0\]

OpenStudy (anonymous):

now from t-9=0 -> t=9

OpenStudy (anonymous):

sub back the t value

OpenStudy (anonymous):

\[3^{x}=9\]

OpenStudy (anonymous):

English plz xD (I dont needa know this stuff I'm just looking at wat u guys r doing)

OpenStudy (anonymous):

\[3^{2}=9\]

OpenStudy (anonymous):

\[3^{x}=3^{2}\]

OpenStudy (anonymous):

and for the other eq 3y-1=0 -> y=-1/3 has no solution

OpenStudy (anonymous):

so x=2

OpenStudy (jadedry):

ah yes of course, I'm so stupid. X.X - I forgot to convert the "t" back to the 3x^2

OpenStudy (anonymous):

thats ok hope it helps :)

OpenStudy (jadedry):

Thank's ayeshaa, I had some other problems with the same line of thought, I'll be able to solve them now. ;u;

OpenStudy (jadedry):

Thank you! I'm closing the board now.